OpenStudy (anonymous):

How long it will take for $300 to double when invested at 6% annual interest compounded twice a year? 10.3 years 11.7 years 12.5 years 13.1 years 13.7 years How long it will take for $300 to double when invested at 6% annual interest compounded twice a year? 10.3 years 11.7 years 12.5 years 13.1 years 13.7 years @Mathematics

OpenStudy (anonymous):

600=300 (1.03)^t 2 = 1.03^t

OpenStudy (anonymous):

ln(2)=t ln(1.03) t=ln(2)/ln(1.03) t= 23.44 about 11.72 years

OpenStudy (anonymous):

Where did the 600 come from?

OpenStudy (anonymous):

start with 300 ... double to get the 600!!!

OpenStudy (anonymous):

Oh i understand now, but how did you know to use ln instead of log